OpenStudy (anonymous):

What is the slope of a line defined by the equation 3x + y = -4?

Nnesha (nnesha):

-3 is a slope

OpenStudy (anonymous):

What is the slope of the line defined by the equation 3x - y = 0

OpenStudy (anonymous):

To find the slope you have to isolate the y\[3x+y=4\]\[y=-3x+4\]Now the slope is the term that multiple the X term\[\boxed{\boxed{slope~=~m~=~-3}}\]

OpenStudy (anonymous):

thanks got it do you understand the next one........please help

OpenStudy (anonymous):

i understand it i just am double checking everything......i have to pass the course to graduate

OpenStudy (anonymous):

is it -3 or 3

OpenStudy (anonymous):

-3 you change the sides, then you change de signal

OpenStudy (anonymous):

What is the slope of the line defined by the equation 2x + 3y = 18

OpenStudy (anonymous):

Now \[3x-y=0\] You isolate the Y again\[y=3x\]therefore the second answer is \[\boxed{\boxed{slope~=~m~=~3}}\]

OpenStudy (anonymous):

you just said it was a negative

OpenStudy (anonymous):

Again isolate the Y\[2x+3y=18\]\[3y=18-2x\]divide each member by 3\[y=6-\frac{2}{3}x\]therefore:\[slope~=~m~=~-\frac{2}{3}\]

OpenStudy (anonymous):

The first answer slope = -3 The second answer slope = 3 The third answer slope = -2/3

OpenStudy (anonymous):

??? what about this

OpenStudy (anonymous):

B i think cause it crosses 2.5

Nnesha (nnesha):

A local utilities department is responsible for billing its town's citizens for their water usage each month. The amount of each household's bill is determined by the number of gallons of water they use monthly. For example, If a household uses between 1000 and 2000 gallons of water, their bill will be $57. However, if they use between 2001 gallons and 3000 gallons, their bill will be $72. Which type of function would best model this situation: a step function or a piecewise function that is not a step function? Explain. can you help me in this question please??

OpenStudy (anonymous):

im not trying to be funny but their is a reason im asking for help on math i have all As but in math.............

OpenStudy (anonymous):

i dont want to give u the wrong answer

Nnesha (nnesha):

okkk that's fine